# The effect of priors on Learning with Restricted Boltzmann Machines

### v1

@@ Abstract
Introduction
Restricted Boltzmann Machines (RBMs) [13, 43, 38, 25] are a type of generative stochastic neural network (NN) that can learn a probability distribution over its set of inputs. Their ability to learn rich internal representations [41, 23, 21, 20, 26] makes them a fundamental building block in deep learning architectures [34, 33, 6, 5, 28, 29].
RBMs have a visible layer of $N$ units $\bm{s}=\left\{s_{i}\right\}_{1\leq i\leq N}$, a hidden layer of $P$ units $\bm{\tau}=\left\{\tau_{\mu}\right\}_{1\leq\mu\leq P}$ and a set of internal connections $\bm{\xi}=\left\{\xi^{\mu}_{i}\right\}_{1\leq i\leq N,1\leq\mu\leq P}$, which are commonly referred to as weights. Given $\bm{\xi}$, the joint distribution of the visible and hidden layers has the Gibbs structure
@@ Introduction
where $\mathbb{E}_{\tau}$ is the expectation over the hidden units. Marginal Gibbs distributions of the form (2) are also known as generalized Hopfield networks [19, 35, 49, 3, 2, 1, 4]. RBMs with generic priors, i.e. the generalized Hopfield networks, have been extensively studied in the statistical mechanics literature, for their properties as models of associative memory [8, 9] and their challenging connection with the Parisi theory of spin glasses [44, 15, 47, 24, 14, 10]. This studies only address the so-called direct problem, where the weights are either fixed or randomly sampled from fixed distributions, and the statistical properties of the generative model are investigated.
In a machine learning context, however, the weights are learned [30, 48, 52] from a dataset of examples, whose distribution the RBM must reconstruct. Therefore the study of the inverse problem of weights optimization is fundamental for a theoretical understanding of the RBM learning mechanisms. From this perspective, a very useful approach is the so called teacher-student setting where a student RBM is trained with data produced by another teacher RBM [7, 39, 27, 37, 18, 36, 50, 42]. Such studies are crucial to isolate individual characteristics of data and machine architecture in a controlled environment and explain their effects on the NN training. For example in [7] the effects of the inference temperature in relation with the dataset’s size and noise on RBM learning is investigated. In [50] the choice of the hidden layer’s size in relation to the number of patterns in the data and their correlation is studied. In [51] the role of possible multi-body interactions is analyzed in the context of the so called Dense Hopfield models.
In this paper, we explore the impact of unit priors on the learning capacity of RBMs and their different learning phases. We extended the results of [7] on the critical dataset’s size necessary for an effective learning by characterizing the entire learning phase diagram. We show the existence of a spin-glass phase where the machine learns spourius patterns of information, an example retrieval phase where the machine tries to learn by memorization and a signal retrieval phase where conversely it learns by generalization. The existence, size and location of these phases strongly depend on the unit priors’ choice.
@@ 1 The Teacher-Student framework
Consider the inference problem of a student machine (S-RBM) trained over a dataset generated by a teacher machine (T-RBM). The T-RBM model is built upon $P$ quenched patterns $\bm{\hat{\xi}}\sim P(\hat{\bm{\xi}})$. A dataset of $M$ examples ${\bm{\mathcal{S}}:=\{\bm{s}^{a}\}^{M}_{a=1}}=\{s^{a}_{1},\ldots,s^{a}_{N}\}^{M}_{a=1}$ is generated by drawing independent samples from the T-RBM distribution
@@
and provided to the student. The student is trained over $\bm{\mathcal{S}}$ to recover its structure, i.e. the teacher patterns. In this process the student patterns $\bm{\xi}$ are optimized. In a Bayesian framework, they are sampled from the posterior distribution
@@
In general, the student is unaware of the properties of the T-RBM, therefore we assume $\hat{\beta}\neq\beta$, $P(\bm{\hat{\xi}})\neq P(\bm{\xi})$, $P(\bm{\hat{\tau}})\neq P(\bm{\tau})$. Without the term $z^{-M}(\bm{\xi})$, the posterior distribution (4) would correspond to $P$ independent generalized Hopfield models, one for each student pattern $\bm{\xi}^{\mu}$, where the examples $\{\bm{s}^{a}\}_{a=1}^{M}$ act as dual patterns [18, 27, 7, 51]. In [50] it is shown that, for teacher patterns that are uncorrelated, the interaction term $z(\bm{\xi})^{-M}$ has the only effect of enforcing orthogonality between student patterns. Furthermore, the learning performance of a S-RBM with $P$ hidden units learning $P$ teacher patterns is equivalent to that of $P$ separate S-RBM with a single hidden unit, each learning one pattern. For this reason, in the rest of the paper we assume $P=1$, see Fig. 1.
@@
where $\Omega_{x}\in[0,1]$, $g_{x}\sim N(0,1)$ and $\epsilon_{x}$ is a Rademacher random variables taking values ±$\pm 1$. We denote its probability distribution with $P_{\Omega_{x}}$, which interpolates between a standard gaussian and a binary distributions, having zero mean and unitary variance for any choice of $\Omega_{x}$. We assume the entries of the teacher pattern $\{\hat{\xi}_{i}\}_{i=1}^{N}$ are drawn independently from $P_{\Omega_{\hat{\xi}}}$. Similarly $\hat{\tau}\sim P_{\Omega_{\hat{\tau}}}$, $\tau\sim P_{\Omega_{\tau}}$, $\xi_{i}\sim P_{\Omega_{\xi}}$ and $s^{\mu}_{i}\sim P_{\Omega_{s}}$. We indicate with $\bm{\Omega}=\{\Omega_{x}\}_{x\in\{\hat{\xi},\tau,\xi,\tau,s\}}$ the set of prior hyperparameters.
@@
$-\beta f(\bm{\Omega};\alpha,\hat{\beta},\beta)=\lim_{M,N\to\infty}\frac{1}{N}\mathbb{E}_{\bm{\hat{\xi}},\bm{\bm{S}}}\log\left[Z(\bm{S})\right],$ (8)
@@ 2 Free energy and saddle point equations
The quenched free energy can be computed exploiting the replica method in the replica symmetry (RS) approximation (see the Appendix A) and reads as
$-\beta f^{RS}(\bm{\Omega};\alpha,\hat{\beta},\beta)=-\beta\operatorname{Extr}_{\Lambda,\Lambda_{\tau},\hat{\Lambda},\hat{\Lambda}_{\tau}}\hat{f}(\Lambda,\Lambda_{\tau},\hat{\Lambda},\hat{\Lambda}_{\tau})\,,$ (9)
where we introduced the sets of parameters $\Lambda=\{p,m,q,d\}$ and $\Lambda_{\tau}=\{p_{\tau},m_{\tau},q_{\tau},d_{\tau}\}$, together with their conjugates $\hat{\Lambda}$ and $\hat{\Lambda}_{\tau}$. The function $\hat{f}$ to be extremized is defined as
$-\beta\hat{f}(\Lambda,\Lambda_{\tau},\hat{\Lambda},\hat{\Lambda}_{\tau})$ $=\frac{1}{2}\hat{q}q+\alpha\frac{1}{2}\hat{q}_{\tau}q_{\tau}-\hat{m}m-\alpha\hat{m}_{\tau}m_{\tau}-\hat{d}d-\alpha\hat{d}_{\tau}d_{\tau}+\alpha^{2}\frac{\beta}{2}\Omega_{\xi}p_{\text{$\tau$}}^{2}+\frac{\beta}{2}\Omega_{\tau}p^{2}+$ (10)
$+\frac{\alpha}{\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}}\Bigg{\langle}\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{s}\log\mathbb{E}_{\tau}\exp\Bigg{\{}-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\hat{p}_{\tau}s)\Bigg{\}}\Bigg{\rangle}_{z}+$
$+\frac{1}{\mathbb{E}_{\hat{\xi}}e^{\hat{d}_{\hat{\xi}}\hat{\xi}^{2}}}\Bigg{\langle}\mathbb{E}_{\hat{\xi}}e^{\hat{d}_{\hat{\xi}}\,\hat{\xi}^{2}}\mathbb{E}_{s}\log\mathbb{E}_{\xi}\exp\Bigg{\{}-(\frac{\hat{q}}{2}-\hat{d})\xi^{2}+\xi(\hat{m}\hat{\xi}+\sqrt{\hat{q}}z+\hat{p}s)\Bigg{\}}\Bigg{\rangle}_{z}+$
$-\frac{\beta\alpha}{2}qq_{\tau}+\alpha\sqrt{\hat{\beta}\beta}mm_{\tau}-\alpha p_{\tau}\hat{p}_{\tau}-p\hat{p}+\frac{\beta\alpha}{2}dd_{\tau}-\alpha\log\mathbb{E}_{\tau}e^{\frac{\beta}{2}d\,\tau^{2}}\,,$
where .$\left\langle.\right\rangle_{z}$ denotes the expectation w.r.t. a standard gaussian random variable $z\sim\mathcal{N}(0,1)$. Extremization gives
$\hat{p}_{\tau}$ $=\beta\alpha\,\Omega_{\xi}\,p_{\tau},\quad\hat{m}_{\tau}=\sqrt{\beta\hat{\beta}}m,\quad\hat{q}_{\tau}=\beta q,\quad\hat{d}_{\tau}=\beta\frac{d}{2},$ (11)
$\hat{p}$ $=\beta\Omega_{\tau}p,\quad\hat{m}=\alpha\sqrt{\beta\hat{\beta}}m_{\tau},\quad\hat{q}=\alpha\beta q_{\tau},\quad\hat{d}=\frac{\alpha\beta}{2}\left(d_{\tau}-\langle\tau^{2}\rangle_{\tau}\right),$ (12)
for the conjugated parameters where we define $\langle\tau^{2}\rangle_{\tau}$ as follows $\langle\tau^{2}\rangle_{\tau}=\mathbb{E}_{\tau}[\tau^{2}e^{\frac{\beta}{2}d\,\tau^{2}}]/\mathbb{E}_{\tau}[e^{\frac{\beta}{2}d\,\tau^{2}}])=(1-\beta\,d\,\Omega_{\tau}^{2})/(1-\beta\,d\,\Omega_{\tau})^{2}$. The order parameters have to be solutions of the following saddle point equations:
@@
The averages .$\left\langle.\right\rangle_{\hat{\xi}}$ and .$\left\langle.\right\rangle_{s}$ are respectively w.r.t. $P_{\Omega_{\hat{\xi}}}(\hat{\xi})$ and $P_{\Omega_{s}}(s)$, while .$\left\langle.\right\rangle_{\hat{\tau}}$ is over $P_{\Omega_{\hat{\tau}}}(\hat{\tau})e^{\frac{\hat{\beta}}{2}\hat{\tau}^{2}}$. Finally .$\left\langle.\right\rangle_{\xi|z,s,\hat{\xi}}$ and .$\left\langle.\right\rangle_{\tau|z,s,\hat{\tau}}$ stand for the expectation over respectively
$P_{\Omega_{\xi}}(\xi)e^{-\frac{\alpha}{2}\beta\left(q_{\tau}-d_{\tau}+\langle\tau^{2}\rangle_{\tau}\right)\xi^{2}+\xi\left(\alpha\sqrt{\hat{\beta}\beta}m_{\tau}\hat{\xi}+\sqrt{\alpha\beta q_{\tau}}z+\beta\Omega_{\tau}ps\right)}\,,$ (17)
$P_{\Omega_{\tau}}(\tau)e^{-\frac{\beta}{2}\left(q-d\right)\tau^{2}+\tau\left(\sqrt{\hat{\beta}\beta}m\hat{\tau}+\sqrt{\beta q}z+\beta\alpha\Omega p_{\tau}s\right)}\,.$ (18)
The optimal order parameters $p,m,q,d$ solving Eqs. (13-16) have to be interpreted as expected overlaps. First of all, $m$ is the limiting expected overlap between the teacher pattern $\bm{\hat{\xi}}$ and the student pattern $\bm{\xi}$, i.e.
@@
The RS saddle-point Eqs.(13-16) can be solved by numerical iteration for any values of the hyper-parameters $\hat{\beta}$, $\beta$, $\alpha$ and $\bm{\Omega}$. We expect the RS solution to be exact when the student is fully informed about the teacher’s prior and hyperparameters and matches them with its own, i.e. $\beta=\hat{\beta}$, $\Omega_{\hat{\xi}}=\Omega_{\xi}$ and $\Omega_{\hat{\tau}}=\Omega_{\tau}$. This regime of complete information is commonly referred to as the Bayes Optimal setting or the Nishimori line [45, 40, 46, 22]. Outside of this regime, i.e. in the mismatched setting, replica symmetry breaking corrections are expected at low temperature. It is important to remark on the role of $z(\bm{\xi})^{-M}$ in Eq. (5). This term is responsible for the emergence of the quadratic term $-\frac{\alpha\beta}{2}\xi^{2}\langle\tau^{2}\rangle_{\tau}$ within the distribution (17). This mechanism enables the automatic regularization of the self-overlap, consistently yielding $d=1$ as the solution to Eq. (16), as will be shown throughout the following sections, and in the Appendix (A).
@@ 3 Exploring arbitrary priors: the Bayesian Optimal setting
With the set of equations (13-16), we can gain deeper insights into what is happening in the student learning process. In this Section we analyze the ideal situation referred to as the Bayesian optimality setting [12], in which the student has access to all the information about the generating process and therefore it is able to mimic the teacher architecture and hyper-parameters to have a better chance of recovering the ground truth $\bm{\hat{\xi}}$. In this regime we therefore assume $\hat{\beta}=\beta$, $\Omega_{\hat{\xi}}=\Omega_{\xi}$ and $\Omega_{\hat{\tau}}=\Omega_{\tau}$. At low temperature all the examples have macroscopic alignment with the signal and the student can easily learn it perfectly. Conversely, at a sufficiently high generating temperature $\hat{\beta}$ the examples have vanishing overlap with the teacher pattern. At the same time, the high inference temperature $\beta=\hat{\beta}$ prevents the student pattern to have macroscopic alignment with a single example, i.e. $p=0$. The saddle point equations (13-16) reduces to
@@
We can further reduce the number of equations thanks to the Nishimori identities [22, 40, 45]. Infact, in the Bayes optimal setting it holds
@@
As a consequence, recalling Eqs. (19,21), it must be $m=q$ and therefore it is sufficient to solve
@@
$P_{\Omega_{\xi}}(\xi)e^{-\frac{\alpha}{2}\beta\left(m_{\tau}-d_{\tau}+\langle\tau^{2}\rangle_{\tau}\right)\xi^{2}+\xi\sqrt{\alpha\beta m_{\tau}}z}\,,$ (30)
$P_{\Omega_{\tau}}(\tau)e^{-\frac{\beta}{2}\left(m-d\right)\tau^{2}+\tau\sqrt{\beta m}z}\,.$ (31)
@@
In the ferromagnetic phase the student RBM is capable of learning (at least partially) the teacher pattern. Conversely this ability is lost in the paramagnetic phase. The ferromagnetic phase emerges even at high temperature once the dataset size reaches a critical threshold, i.e. $\alpha\geq\alpha_{c}(T,\bm{\Omega})$. In fact, at high generating temperature, each example provides only a vanishing amount of information about the teacher pattern. Therefore an extensively large number of examples are needed for the student to retrieve the signal. The corresponding P-F phase transition is thus the onset of learning. Descending from high temperatures, where $m=m_{\tau}=0$, the P-F transition line can be obtained by expanding in $m\sim 0$. First of all, note that at $m=m_{\tau}=0$ Eqs. (28, 29) give
$d_{\tau}=\langle\tau^{2}\rangle_{0}=\mathbb{E}_{\tau}[\tau^{2}e^{\frac{\beta}{2}d\,\tau^{2}}]/\mathbb{E}_{\tau}[e^{\frac{\beta}{2}d\,\tau^{2}}]\,,\qquad\tau\sim P_{\Omega_{\tau}}$ (32)
$d=\langle\xi^{2}\rangle_{0}=\mathbb{E}_{\xi}[\xi^{2}e^{\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\,\xi^{2}}]/\mathbb{E}_{\xi}[e^{\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\,\xi^{2}}]\,,\ \ \ \ \ \ \xi\sim P_{\Omega_{\xi}}$ (33)
where $\langle\cdot\rangle_{0}$ denotes the expectations w.r.t. the distributions of Eqs. (30) or (31) at zero effective field. Note that, since $d_{\tau}=\langle\tau^{2}\rangle_{\tau}$ it is $e^{\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\,\xi^{2}}=1$ and therefore $d=1$. This is expected because, thanks to the Nishimori identities, it must be
$\mathbb{E}_{\bm{\hat{\xi}},\bm{S},\bm{\xi}}f(\bm{\xi})=\,\mathbb{E}_{\bm{\hat{\xi}},\bm{S},\bm{\xi}}f(\bm{\hat{\xi}})=\mathbb{E}_{\bm{\hat{\xi}}}f(\bm{\hat{\xi}}).$ (34)
Therefore the statistics of the student pattern matches those of the teacher’s: in particular the self-overlap must be equal to that of the teacher pattern, being one by definition. By expanding the effective distributions in Eqs. (30, 31) as
@@
$P_{\Omega_{\tau}}(\tau)\left(1+\sqrt{\beta m}z\,\tau+O(q)\right)e^{\frac{\beta}{2}\tau^{2}}\,$ (36)

### v2

@@ Abstract
keywords:
Statistical Mechanics , Machine Learning , self-supervised learning
journal: Physica A
[…]
[label1]organization=Department of Mathematics, University of Bologna, addressline=Piazza di Porta San Donato 5, city=Bologna (BO), postcode=40126, country=Italy
{highlights}
A teacher-student unsupervised learning problem analyzed with Restricted Boltzmann Machines with general priors.
Phase diagrams, derived in the statistical mechanics perspective, reveal different learning regimes and are sensitive to the choice of priors, ranging from gaussian to binary.
A critical training set size can be defined as a triple point in the phase diagram, below which learning by generalization becomes impossible.
Hidden unit priors with Gaussian tails facilitate learning by reducing the critical size or the level of regularization required to enter the generalization phase.
1 Introduction
Restricted Boltzmann Machines (RBMs) barra2012equivalence , mehta2019high , huang2022statistical , decelle2021restricted are a type of generative stochastic neural network (NN) that can learn a probability distribution over its set of inputs. Their ability to learn rich internal representations le2008representational , decelle2017spectral , carbone2024fast , bereux2024fast , decelle2024inferring makes them a fundamental building block in deep learning architectures RBMs have a visible layer of $N$ units $\bm{s}=\left\{s_{i}\right\}_{1\leq i\leq N}$, a hidden layer of $P$ units $\bm{\tau}=\left\{\tau_{\mu}\right\}_{1\leq\mu\leq P}$ and a set of internal connections $\bm{\xi}=\left\{\xi^{\mu}_{i}\right\}_{1\leq i\leq N,1\leq\mu\leq P}$, which are commonly referred to as weights. Given $\bm{\xi}$, the joint distribution of the visible and hidden layers has the Gibbs structure
@@ 1 Introduction
where $\mathbb{E}_{\tau}$ is the expectation over the hidden units. Marginal Gibbs distributions of the form (2) are also known as generalized Hopfield networks barra2018phase , hopfield1982neural , sollich2014extensive , agliari2017neural , agliari2015retrieval , agliari2015hierarchical , agliari2018non . RBMs with generic priors, i.e. the generalized Hopfield networks, have been extensively studied in the statistical mechanics literature, for their properties as models of associative memory amit1985storing , amit1987statistical and their challenging connection with the Parisi theory of spin glasses mezard2009information , barra2015multi , panchenko2015free , decelle2018thermodynamics , barra2023thermodynamics , antenucci2019approximate . This studies only address the so-called direct problem, where the weights are either fixed or randomly sampled from fixed distributions, and the statistical properties of the generative model are investigated.
In a machine learning context, however, the weights are learned fischer2014training , salakhutdinov2007restricted , tubiana2018restricted from a dataset of examples, whose distribution the RBM must reconstruct. Therefore the study of the inverse problem of weights optimization is fundamental for a theoretical understanding of the RBM learning mechanisms. From this perspective, a very useful approach is the so called teacher-student setting where a student RBM is trained with data produced by another teacher RBM alemanno2023hopfield , huang2016unsupervised , decelle2021inverse , huang2017statistical , barra2017phase , hou2019minimal , theriault2024modelling , lesieur2017constrained . Such studies are crucial to isolate individual characteristics of data and machine architecture in a controlled environment and explain their effects on the NN training. For example in alemanno2023hopfield the effects of the inference temperature in relation with the dataset’s size and noise on RBM learning is investigated. In theriault2024modelling the choice of the hidden layer’s size in relation to the number of patterns in the data and their correlation is studied. In theriault2024dense , agliari2023denseunsupervised , agliari2023densesupervised , the role of multi-body interactions is analyzed in the context of Dense Hebbian neural networks.
In this paper, we explore the impact of unit priors on the learning capacity of RBMs and their different learning phases. We extended the results of alemanno2023hopfield on the critical dataset’s size necessary for an effective learning by characterizing the entire learning phase diagram. We show the existence of a spin-glass phase where the machine learns spourius patterns of information, an example retrieval phase where the machine tries to learn by memorization and a signal retrieval phase where conversely it learns by generalization. The existence, size and location of these phases strongly depend on the unit priors’ choice.
@@ 2 The Teacher-Student framework
[…]
@@
and provided to the student.
The student is trained over $\bm{\mathcal{S}}$ to recover its structure, i.e. the teacher patterns. In this process the student patterns $\bm{\xi}$ are optimized. In a Bayesian framework, they are sampled from the posterior distribution
@@
In general, the student is unaware of the properties of the T-RBM, therefore we assume $\hat{\beta}\neq\beta$, $P(\bm{\hat{\xi}})\neq P(\bm{\xi})$, $P(\bm{\hat{\tau}})\neq P(\bm{\tau})$. Without the term $z^{-M}(\bm{\xi})$, the posterior distribution (4) would correspond to $P$ independent generalized Hopfield models, one for each student pattern $\bm{\xi}^{\mu}$, where the examples $\{\bm{s}^{a}\}_{a=1}^{M}$ act as dual patterns barra2017phase , decelle2021inverse , alemanno2023hopfield , theriault2024dense . In theriault2024modelling it is shown that, for teacher patterns that are uncorrelated, the interaction term $z(\bm{\xi})^{-M}$ has the only effect of enforcing orthogonality between student patterns. Furthermore, the learning performance of a S-RBM with $P$ hidden units learning $P$ teacher patterns is equivalent to that of $P$ separate S-RBM with a single hidden unit, each learning one pattern. For this reason, in the rest of the paper we assume $P=1$, see Fig. 1.
@@
where $\Omega_{x}\in[0,1]$, $g_{x}\sim N(0,1)$ and $\epsilon_{x}$ is a Rademacher random variables taking values $\pm 1$. We denote its probability distribution with $P_{\Omega_{x}}$, which interpolates between a standard gaussian and a binary distributions, having zero mean and unitary variance for any choice of $\Omega_{x}$. We assume the entries of the teacher pattern $\{\hat{\xi}_{i}\}_{i=1}^{N}$ are drawn independently from $P_{\Omega_{\hat{\xi}}}$. Similarly $\hat{\tau}\sim P_{\Omega_{\hat{\tau}}}$, $\tau\sim P_{\Omega_{\tau}}$, $\xi_{i}\sim P_{\Omega_{\xi}}$ and $s^{\mu}_{i}\sim P_{\Omega_{s}}$. We indicate with $\bm{\Omega}=\{\Omega_{x}\}_{x\in\{\hat{\xi},\tau,\xi,\tau,s\}}$ the set of prior hyperparameters.
@@
$\beta f(\bm{\Omega};\alpha,\hat{\beta},\beta)=\lim_{M,N\to\infty}\frac{1}{N}\mathbb{E}_{\bm{\hat{\xi}},\bm{\bm{S}}}\log\left[Z(\bm{S})\right],$ (8)
@@ 3 Free energy and saddle point equations
The quenched free energy can be computed by exploiting the replica method in the replica symmetry (RS) approximation and reads as
$\beta f^{RS}(\bm{\Omega};\alpha,\hat{\beta},\beta)=-\beta\operatorname{Extr}_{\Lambda,\Lambda_{\tau},\hat{\Lambda},\hat{\Lambda}_{\tau}}\hat{f}(\Lambda,\Lambda_{\tau},\hat{\Lambda},\hat{\Lambda}_{\tau})\,$, (9)
where we introduced the sets of parameters $\Lambda=\{p,m,q,d\}$ and $\Lambda_{\tau}=\{p_{\tau},m_{\tau},q_{\tau},d_{\tau}\}$, together with their conjugates $\hat{\Lambda}$ and $\hat{\Lambda}_{\tau}$. The function $\hat{f}$, together with its full derivation, can be found in A, specifically in Eq.(89). It is extremized by solving the following equations for the order parameters
@@
[…]
$P_{\Omega_{\xi}}(\xi)\exp\left({-\frac{\alpha}{2}\beta\left(q_{\tau}-d_{\tau}+\langle\tau^{2}\rangle_{\tau}\right)\xi^{2}+\xi\left(\alpha\sqrt{\hat{\beta}\beta}m_{\tau}\hat{\xi}+\sqrt{\alpha\beta q_{\tau}}z+\beta\Omega_{\tau}ps\right)}\right)\,$, (14)
$P_{\Omega_{\tau}}(\tau)\exp{\left(-\frac{\beta}{2}\left(q-d\right)\tau^{2}+\tau\left(\sqrt{\hat{\beta}\beta}m\hat{\tau}+\sqrt{\beta q}z+\beta\alpha\Omega p_{\tau}s\right)\right)}\,.$ (15)
The optimal order parameters $p,m,q,d$ solving Eqs. (10-13) have to be interpreted as expected overlaps. First of all, $m$ is the limiting expected overlap between the teacher pattern $\bm{\hat{\xi}}$ and the student pattern $\bm{\xi}$, i.e.
@@
The RS saddle-point Eqs.(10-13) can be solved by numerical iteration for any values of the hyper-parameters $\hat{\beta}$, $\beta$, $\alpha$ and $\bm{\Omega}$. We expect the RS solution to be exact when the student is fully informed about the teacher’s prior and hyperparameters and matches them with its own, i.e. $\beta=\hat{\beta}$, $\Omega_{\hat{\xi}}=\Omega_{\xi}$ and $\Omega_{\hat{\tau}}=\Omega_{\tau}$. This regime of complete information is commonly referred to as the Bayes Optimal setting or the Nishimori line nishimori1980exact , iba1999nishimori , nishimori2001statistical , contucci2009spin .
Outside of this regime, i.e. in the mismatched setting, replica symmetry breaking corrections are expected at low temperature. It is important to remark on the role of $z(\bm{\xi})^{-M}$ in Eq. (5). This term is responsible for the emergence of the quadratic part $\frac{\alpha\beta}{2}\xi^{2}\langle\tau^{2}\rangle_{\tau}$ within the distribution (14). This mechanism enables the automatic regularization of the self-overlap, consistently yielding $d=1$ as the solution to Eq. (13), as will be shown throughout the following sections, and in A.
@@ 4 Exploring arbitrary priors: the Bayesian Optimal setting
With the set of equations (10-13), we can gain deeper insights into what is happening in the student learning process. In this Section we analyze the ideal situation referred to as the Bayesian optimality setting barbier2022bayes , in which the student has access to all the information about the generating process and therefore it is able to mimic the teacher architecture and hyper-parameters to have a better chance of recovering the ground truth $\bm{\hat{\xi}}$. In this regime we therefore assume $\hat{\beta}=\beta$, $\Omega_{\hat{\xi}}=\Omega_{\xi}$ and $\Omega_{\hat{\tau}}=\Omega_{\tau}$. At low temperature all the examples have macroscopic alignment with the signal and the student can easily learn it perfectly. Conversely, at a sufficiently high generating temperature $\hat{\beta}$ the examples have vanishing overlap with the teacher pattern. At the same time, the high inference temperature $\beta=\hat{\beta}$ prevents the student pattern to have macroscopic alignment with a single example, i.e. $p=0$. The saddle point equations (10-13) reduces to
@@
We can further reduce the number of equations thanks to the Nishimori identities contucci2009spin , iba1999nishimori , nishimori1980exact . Infact, in the Bayes optimal setting it holds
@@
As a consequence, recalling Eqs. (16,18), it must be $m=q$ and therefore it is sufficient to solve
@@
$P_{\Omega_{\xi}}(\xi)\exp{\left(-\frac{\alpha}{2}\beta\left(m_{\tau}-d_{\tau}+\langle\tau^{2}\rangle_{\tau}\right)\xi^{2}+\xi\sqrt{\alpha\beta m_{\tau}}z\right)}\,$, (27)
$P_{\Omega_{\tau}}(\tau)\exp{\left(-\frac{\beta}{2}\left(m-d\right)\tau^{2}+\tau\sqrt{\beta m}z\right)}\,.$ (28)
@@
In the ferromagnetic phase the student RBM is capable of learning (at least partially) the teacher pattern. Conversely this ability is lost in the paramagnetic phase. The ferromagnetic phase emerges even at high temperature once the dataset size reaches a critical threshold, i.e. $\alpha\geq\alpha_{c}(T,\bm{\Omega})$. In fact, at high generating temperature, each example provides only a vanishing amount of information about the teacher pattern. Therefore an extensively large number of examples are needed for the student to retrieve the signal. The corresponding P-F phase transition is thus the onset of learning. Descending from high temperatures, where $m=m_{\tau}=0$, the P-F transition line can be obtained by expanding in $m\sim 0$. First of all, note that at $m=m_{\tau}=0$ Eqs. (25, 26) give
$d_{\tau}=\langle\tau^{2}\rangle_{0}=\mathbb{E}_{\tau}[\tau^{2}\exp{\left(\frac{\beta}{2}d\,\tau^{2}\right)}]/\mathbb{E}_{\tau}[\exp{\left(\frac{\beta}{2}d\,\tau^{2}\right)}]\,,\qquad\tau\sim P_{\Omega_{\tau}}$ (29)
$\hskip-62.59596ptd=\langle\xi^{2}\rangle_{0}=\mathbb{E}_{\xi}[\xi^{2}\exp{\left(\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\,\xi^{2}\right)}]/\mathbb{E}_{\xi}[\exp{\left(\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\,\xi^{2}\right)}]\,,\ \ \xi\sim P_{\Omega_{\xi}}$ (30)
where $\langle\cdot\rangle_{0}$ denotes the expectations w.r.t. the distributions of Eqs. (27) or (28) at zero effective field. Note that, since $d_{\tau}=\langle\tau^{2}\rangle_{\tau}$ it is $\exp\left(\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\,\xi^{2}\right)=1$ and therefore $d=1$. This is expected because, thanks to the Nishimori identities, it must be
[…]
Therefore the statistics of the student pattern matches those of the teacher’s: in particular the self-overlap must be equal to that of the teacher pattern, being one by definition. By expanding the effective distributions in Eqs. (27, 28) as
@@
$P_{\Omega_{\tau}}(\tau)\left(1+\sqrt{\beta m}z\,\tau+O(q)\right)\exp{\left\{\frac{\beta}{2}\tau^{2}\right\}}\,$ (33)